How many no atoms are present in 0.1 mole of of p4 atom?
 

Grade:11

1 Answers

Arun
6 years ago
Number of molecules = number of moles × Avogadro's constant .
no of molecules = 0.1 × 6.023 × 10²³ 
= 6.023 × 10²² 
in P4 
no of atoms = 4 × no of molecules 
so, 
no of atoms = 4 × 6.023 × 10²²
= 24.092 × 10²²
= 2.4092 × 10²³